40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 248 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  4483   Fri Apr 1 23:49:24 2011 kiwamuSummaryGreen Lockingtwo states in green beat-note

According to the measurement done by Aidan and me, there are two beat-note state.

One gave us a small beat signal and the other gave us a bigger signal by approximately 20 dB.

 

 A possible reason for this phenomenon is that the end laser is operating at a special temperature that somehow drives the laser with two different modes at the same time.

So that it permits the laser sometimes locked with one of the two modes and sometimes with the other mode.

For the first step we will change the temperature such that the laser can run with a single stable mode.

Then for investigating it we will put a scanning cavity on the X end table to see if it really exhibits a two modes or not.

Quote from #4472

The attached table shows the amplitude of the green beat note when the end laser was in various states. We can increase the beat note amplitude dramatically by switching to a different states.

  4482   Fri Apr 1 23:05:58 2011 kiwamuUpdateGreen Lockingnoise budget

I made a coarse noise budget in order to decide our next actions for the X arm green locking.

So be careful, this is not an accurate noise budget !

 Some data are just coming from rough estimations and some data are not well calibrated.

noise_budget.png

 Assuming all the noise are not so terribly off from the true values, the noise at high frequency is limited by the dark noise of the PD or it already reaches to the IR inloop signal.

The noise at low frequency is dominated by the intensity noise from the transmitted green light although we thought it has been eliminated by the comparator.

In any case I will gradually make this noise budget more accurate  by collecting some data and calibrating them.

 

According to the plot what we should do are :

  * More accurate PD noise measurement

  * More accurate shot noise estimation

  * Searching for a cause of the small beat signal (see here) because a bigger beat signal lowers the PD noise.

  * Investigation of the Intensity noise

  4481   Fri Apr 1 18:54:41 2011 BryanConfigurationGreen LockingY end doubling oven

The doubling oven is now ready to go for the Y arm. The PPKTP crystal is mounted in the oven:

P4010036.JPG

Note - the crystal isn't as badly misaligned as it looks in this photo. It's just an odd perspective shot. I then closed it up and checked to make sure the IR beam on the Y bench passes through the crystal. It does. Just need to tweak the waist size/position a bit and then we can actually double some frequencies!

P4010041.JPG

  4480   Thu Mar 31 20:46:11 2011 AidanSummaryGreen LockingGreen beat note PD DC response

I measured the DC response of the Green PD


Power into PD at DC (green laser pointer) = 285 uW
Voltage out of PD = 552mV/(100x SR560gain) = 5.52mV
Photocurrent = 5.52mV/(241 Ohms)*3 = 68.7uA
Responsivity = 68.7/285 = 0.24 A/W

Therefore, since the responsivity is in the correct range for a Silicon PD at 532nm, the DC output is giving us sensible response to an input signal.


But, there is a 2.12MHz, 328mV oscillation on the DC output irrespective of the incident power.
 

  4479   Thu Mar 31 20:37:10 2011 AidanSummaryGreen LockingRF amplitude source

 I gutted one of the $2 red laser pointers to build a laser source whose amplitude we could modulate at RF frequencies. Basically, I cut off the bulk of the housing from the pointer and soldered a BNC connection into the two terminals that the 2x 1.5V batteries were connected to. When I applied 3V across this BNC connector the diode still worked. So far so good.

Next I added a bias tee to the input. I put 3V across the DC input of the bias tee and added a -3dBm signal into the RF port of the tee. The laser beam was incident on a PDA100A (bandwidth of 1.7MHz) and, sure enough, Kiwamu and I could see a flat response in the amplitude at a given drive frequency out to around 1.7MHz.

We should check the response on a faster PD to see how fast the laser diode is, but we should be able to use this now to check the RF response of the green beat note PD. 

TO DO:

1. Add some capacitors across the DC input of the bias tee.

2. Do something about the switch on the laser diode.

3. Attach some labels to the laser that specify what is the required DC voltage and the maximum acceptable RF modulation amplitude.

Attachment 1: P1000543.jpg
P1000543.jpg
Attachment 2: P1000544.jpg
P1000544.jpg
Attachment 3: P1000545.jpg
P1000545.jpg
  4478   Thu Mar 31 19:58:11 2011 kiwamuUpdateCDSc1iscex crashed

After I did several things to add new DAQ channels on c1iscex it suddenly became out of network. Maybe crashed.

Then c1iscex didn't respond to a ping and all the epics values associated with c1iscex became not accessible.

I physically shut it down by pushing the reset button. Then it came back and is now running fine.

 


(how I broke it)

Since activateDAQ.py has screwed up the 'ini' files including C1SCX.ini, I was not able to add a channel to C1SCX.ini by the usual daqconfig GUI.

So I started editing it in a manual way with an editor and changed some sentences to that shown below

  [C1:GCX-ERR_MON_IN1_DAQ]
  acquire=1
  chnnum=10004
  datarate=2048
  datatype=4
  [C1:GCX-GRN_REFL_DC_IN1_DAQ]
  acquire=1
  chnnum=10007
  datarate=2048
  datatype=4
  [C1:GCX-SLOW_SERVO1_IN1_DAQ]
  acquire=1
  chnnum=10010
  datarate=2048
  datatype=4

Then I rebooted fb to reflect the new DAQ channels.

After that I looked at the C1_FE_STATUS.adl screen and found some indicator lights were red.

So I pushed "Diag reset" button and "DAQ Reload" button on the C1SCX_GDS_TP.adl screen and then c1iscex died.

After the reboot the new DAQ channels looked acquired happily.

This is my second time to crash a front end machine (see this entry)

  4477   Thu Mar 31 15:23:14 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Quote:

3. Zoomed-in plot for the SHG crystal show no astigmatism. However, the zoomed out plot shows some astigmatism.

How consistent are they? ==> Interested in seeing the fit including the zoomed out measurements.

Right. Fitting to the data. Zoomed out plots first. I used the general equation f(x) = w_o.*sqrt(1 + (((x-z_o)*1064e-9)./(pi*w_o.^2)).^2)+c for each fit which is basically just the Gaussian beam width parameter calculation but with an extra offset parameter 'c'

Vertical fit for zoomed out data:

Coefficients (with 95% confidence bounds):

       c =   7.542e-06  (5.161e-06, 9.923e-06)

       w_o =   3.831e-05  (3.797e-05, 3.866e-05)

       z_o =       1.045  (1.045, 1.046)

 

Goodness of fit:

  SSE: 1.236e-09

  R-square: 0.9994

 
Horizontal fit for zoomed out data:
 

Coefficients (with 95% confidence bounds):

       c =   1.083e-05  (9.701e-06, 1.195e-05)

       w_o =   4.523e-05  (4.5e-05, 4.546e-05)

       z_o =       1.046  (1.046, 1.046)

 

Goodness of fit:

  SSE: 2.884e-10

  R-square: 0.9998

  Adjusted R-square: 0.9998

  RMSE: 2.956e-06

 

Zoomed_out_fitting01.png

-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-

 

OK. Looking at the plots and residuals for this, the deviation of the fit around the waist position, and in fact all over, looks to be of the order 10um. A bit large but is it real? Both w_o values are a bit lower than the 50um we'd like, but… let's check using only the zoomed in data -  hopefully more consistent since it was all taken with the same power setting.

 

 

Vertical data fit using only the zoomed in data:

 

Coefficients (with 95% confidence bounds):

       c =   1.023e-05  (9.487e-06, 1.098e-05)

       w_o =   4.313e-05  (4.252e-05, 4.374e-05)

       z_o =       1.046  (1.046, 1.046)

 

Goodness of fit:

  SSE: 9.583e-11

  R-square: 0.997

 

Horizontal data fit using only the zoomed in data:

 

Coefficients (with 95% confidence bounds):

       c =   1.031e-05  (9.418e-06, 1.121e-05)

       w_o =    4.41e-05  (4.332e-05, 4.489e-05)

       z_o =       1.046  (1.046, 1.046)

 

Goodness of fit:

  SSE: 1.434e-10

  R-square: 0.9951

 

-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-

Zoomed_in_fitting01.png

 

The waists are both fairly similar this time 43.13um and 44.1um and the offsets are similar too  - residuals are only spread by about 4um this time.

 

I'm inclined to trust the zoomed in measurement more due to the fact that all the data was obtained under the same conditions, but either way, the fitted waist is a bit smaller than the 50um we'd like to see. Think it's worthwhile moving the 62.9mm lens back along the bench by about 3/4 -> 1cm to increase the waist size.

 

 

 

 

 

  4476   Thu Mar 31 14:10:00 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Quote:

 I went through the entries.

1. Give us a photo of the day. i.e. Faraday, tilted lens, etc...

2. After all, where did you put the faraday in the plot of the entry 4466?

3. Zoomed-in plot for the SHG crystal show no astigmatism. However, the zoomed out plot shows some astigmatism.
How consistent are they? ==> Interested in seeing the fit including the zoomed out measurements.

 OK. Taking these completely out of order in the easiest first...

2. The FI is between positions 27.75 and 32 on the bench - i.e. this is where the input and output apertures are. (corresponds to between 0.58 and 0.46 on the scale of those two plotsand just before both the vertical and horizontal waists) At these points the beam radius is around 400um and below, and the aperture of the Faraday is 4.8mm (diameter).

1. Photos...

Laser set up - note the odd angles of the mirrors. This is where we're losing a goodly chunk of the light. If need be we could set it up with an extra mirror and send the light round a square to provide alignment control AND reduce optical power loss...

P3310028.JPG

 

Faraday and angled lens - note that the lens angle is close to 45 degrees. In principle this could be replaced with an appropriate cylindrical lens, but as long as there's enough light passing through to the oven I think we're OK.

P3310029.JPG

3. Fitting... coming soon once I work out what it's actually telling me. Though I hasten to point out that the latter points were taken with a different laser power setting and might well be larger than the actual beam width which would lead to astigmatic behaviour.

  4475   Thu Mar 31 11:30:51 2011 steveConfigurationGeneralstrain relieved rf cables

I strain relieved RF cables labeled 33 MHZ LO and 166 MHZ to EOM at 1X2  This is a temporary setup for the 11 MHZ

The coax N  bulkheads connectors are mounted on the plastic front panel now.

Attachment 1: P1070491.JPG
P1070491.JPG
  4474   Thu Mar 31 08:31:44 2011 SureshUpdateVIDEOCable laying...continued

The video work has crossed a milestone.    

Kiwamu and Steve have shifted the three quads from the control room to the Video MUX rack (1Y1) and have wired them to the MUX.

The monitors in the control room have been repositioned and renumbered.  They are now connected directly to the MUX. 

Please see the new cable list for the input and output channels on the MUX.

As of today, all cables according the new plan are in place.  Their status   indicated on the wiki page above is not verified .  Please ignore that column for now, we will be updating that soon.

I shifted the MC1F/MC3F camera and the MC2F cameras onto the new cables.  Also connected the monitors at the BS chamber and end of the X arm to their respective cables.  I have removed the RG58B BNC (black) cables running from MC2 to BS and from ETMXF to the top of the Flow Bench. 

Some of the old video cables are still in place but are not used.   We might consider removing them to clear up the clutter. 

Some of the video cables in use are orange and if the lab's  cable color code is to be enforced these will have to be replaced with blue ones..

Some of the cables in use running from the MUX to the monitor in the control room are the white 50 Ohm variety.  There are also black RG59 Cables running the same way ( we have surplus cables in that path)  and we have to use those instead of the white ones. 

There are a number of tasks remaining:

a)  The inputs from the various existing cameras have to be verified. 

b) There are quite a few cameras which are yet to be installed.

c) The Outputs may not not be connected to their monitors.  That the monitors may still be connected to an old cable which is not connected to the MUX.  The new cable should be lying around close by.  So if you see a blank monitor please connect it to its new cable. 

d) The status column on the wiki page has to be updated.

e) Some of the currently in place may need to be replaced and some need to be removed.  We need to discuss our priorities and come up with a plan for that.

After checking everything we can certify that the video cabling system is complete.

I would like Joon Ho to take care of this verification+documenting process and declaring that the job is complete. 

 

Steve attached these two pictures.

Attachment 1: P1070489.JPG
P1070489.JPG
Attachment 2: P1070494.JPG
P1070494.JPG
  4473   Thu Mar 31 02:59:49 2011 KojiConfigurationGreen LockingThe wonderful world of mode-matching

 I went through the entries.

1. Give us a photo of the day. i.e. Faraday, tilted lens, etc...

2. After all, where did you put the faraday in the plot of the entry 4466?

3. Zoomed-in plot for the SHG crystal show no astigmatism. However, the zoomed out plot shows some astigmatism.
How consistent are they? ==> Interested in seeing the fit including the zoomed out measurements.

  4472   Wed Mar 30 21:46:10 2011 Aidan, KiwamuUpdateGreen LockingStates of the Green beat note

The attached table shows the amplitude of the green beat note when the end laser was in various states. We can increase the beat note amplitude dramatically by switching to a different states.

State 1
C1:GCX-GRN_REFL_DC:             638 counts
C1:GCV-XARM_BEAT_DC: (PSL blocked)    950 avg counts (zero = -794 counts)
amplitude of beat note:            -23dBm (after PD + amps) (f ~ 30 MHz)?
C1:GCX-SLOW_SERVO2_OUT:            318 counts

State 2
C1:GCX-GRN_REFL_DC:             180 counts
C1:GCV-XARM_BEAT_DC: (PSL blocked)    1270 avg counts (zero = -794 counts)
C1:GCV-XARM_BEAT_DC: (PSL unblocked)    1700 avg counts (zero = -794 counts)
amplitude of beat note:            -7dBm (after PD + amps) f = 60MHz
amplitude of beat note:            0dBm (after PD + amps) f = 30MHz
C1:GCX-SLOW_SERVO2_OUT:            290 counts

State 3
C1:GCX-GRN_REFL_DC:             220 counts
C1:GCV-XARM_BEAT_DC: (PSL blocked)    1120 avg counts (zero = -794 counts)
C1:GCV-XARM_BEAT_DC: (PSL unblocked)    1520 avg counts (zero = -794 counts)
amplitude of beat note:            0dBm (after PD + amps) f = 15MHz
C1:GCX-SLOW_SERVO2_OUT:            305 counts

PSL temp = ??
C1:PSL-FSS_SLOWM = -3.524
 

  4471   Wed Mar 30 21:43:31 2011 Aidan, KiwamuSummaryGreen LockingCalculation of the green contrast on the RF PD

Skip to final thought ...

Kiwamu and I have set about measuring the contrast of the signal on the RF PD. We can only do this when the end green laser is locked to the cavity. This is because the green transmission through the cavity, when unlocked, is too low. Unfortunately, once we lock the green beam to the cavity, we can't keep the beatnote on the RF PD stable to within a few hundred Hz of DC (remember that the cavity is swinging around by a couple of FSRs). So we also lock the PSL to cavity.

At this point we're stuck because we can't get both of these beams resonant within the cavity AND have the frequency difference between them be less 1kHz - when the lasers are locked to the cavity, their frequencies are separated by an integer number of FSRs + a fixed frequency offset, f_offset, that is set by the phase difference on reflection from the coating between the two wavelengths (532nm and 1064nm). We can never get the frequency difference between the lasers to be less than this offset frequency AND still have them both locked to the cavity.

 

So our contrast measuring method will have to use the RF signal.

 

So this is our method. We know the incident power from each beam on the RF PD (see Kiwamu's elog entry here), but to recap,

P_green_PSL = 72 uW (as measured today)

P_green_XARM = 560 uW (as measured by Kiwamu last week).

The trans-impedance of the RF PD is 240 Ohms. We'll assume a responsitivity of 0.25 A/W. So, if the XARM transmission and PSL green beams are perfectly matched then the maximum value of the RF beat note should be:

RF_amplitude_max = 2* SQRT(P_green_PSL*P_green_XARM) * responsivity * transimpedance = 240*0.25*2*(72E-6*560E-6)^(1/2) (volts)

= 24 mV = -19.5 dBm (or 27.5dBm after the +47 dB from the two  ZFL-1000LN+ amplifiers - with +15V in - that protrude from the top of the PD)

The maximum RF strength of the beat-note that we measure is around -75 dBm (at the RF output of the PD). This means the contrast is down nearly 600x from optimal. Or it means something is broken.

Final thought: at the end of this procedure we found that the RF beat note amplitude would jump to a different and much higher amplitude state. This renders a lot of the above useless until we discover the cause.

  4470   Wed Mar 30 21:21:15 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Step 4: Matching into the oven

 

 

Now that the astigmatism is substantially reduced, we can work out a lens solution to obtain a 50um waist *anywhere* on the bench as long as there's enough room to work with the beam afterwards. The waist after the Faraday and lens is at position 22.5 on the bench. A 50 mm lens placed 18 cm after this position (position 14.92 on the bench) should give a waist of 50 um at  24.57 cm after the waist (position 12.83 on the bench). This doesn't give much room to measure the beam waist in though - the Beamscan head has a fairly large finite size… wonder if there's a slightly less strong lens I could use…

OK. With a 66 mm lens at 23 cm (position 13.45 on the bench) after the waist we get a 50 um waist at 31.37 cm after the waist (position 10.15 on the bench). 

 

Oven_Lens_Solution_66mm.png

 

Closest lens I found was 62.9mm which will put the 50um point a bit further towards the wall, but on the X-arm the oven is at position 8.75 ish. So anything around there is fine.

 

Using this lens and after a bit of manual fiddling and checking with the Beamscan, I figured we needed a close in, fine-grained measurement so set the Beamscan head up on a micrometer stage Took a whoie bunch of data around position 9 on the bench:

 

 

Position A1_13.5%_width A2_13.5%_width

(mm) (um mean) (um mean)

-15 226.8 221.9

-14 210.9 208.3

-13 195.5 196.7

-12 181.0 183.2

-11 166.0 168.4

-10 154.0 153.1

-9 139.5 141.0

-8 127.5 130.0

-7 118.0 121.7

-6 110.2 111.6

-5 105.0 104.8

-4 103.1 103.0

-3 105.2 104.7

-2 110.9 110.8

-1 116.8 117.0

0 125.6 125.6

0 125.6 125.1

1 134.8 135.3

2 145.1 145.6

3 155.7 157.2

4 168.0 168.1

5 180.5 180.6

6 197.7 198.6

7 211.4 209.7

8 224.0 222.7

9 238.5 233.7

10 250.9 245.8

11 261.5 256.4

12 274.0 270.4

13 291.3 283.6

14 304.2 296.5

15 317.9 309.5

 

Matching_Into_Green_Oven_zoomed_out.pngMatching_Into_Green_Oven_zoomed_in.png

 

And at this point the maximum power available at the oven-waist is 298mW. With 663mW available from the laser with a desired power setting of 700mW on the supply. Should make sure we understand where the power is being lost. The beam coming through the FI looks clean and unclipped, but there is some stray light around.

 

Position A1_13.5%_width A2_13.5%_width

(bench) (um mean) (um mean)

7 868.5   739.9

6 1324 1130

5 1765 1492

4 2214 1862

 

The plot looks pretty good, but again, there looks to be an offset on the 'fitted' curve. Taking a couple of additional points further on to make sure it all works out as the beam propagates. I took a few extra points at the suggestion of Kiwamu and Koji - see the zoomed out plot.  The zoomed in plot has by-eye fit lines - again, because to get the right shape to fit the points there appears to be an offset. Where is that coming from? My suspicion is that the Beamscan doesn't take account of the any background zero offsets when calculating the 13.5% and we've been using low power when doing these measurements - very small focussed beams and didn't want to risk damage to the profiler head.

 

Decided to take a few measurements to test this theory. Trying different power settings and seeing if it gives different offset and/or a changed width size

 

7 984.9 824.0 very low power

7 931.9 730.3 low power

7 821.6 730.6 higher power

7 816.4 729.5 as high as I'm comfortable going

 

Trying this near the waist…

 

8.75 130.09 132.04 low power

8.75 106.58 105.46 higher power

8.75 102.44 103.20 as high as it can go without making it's saturated

 

So it looks like offset *is* significant and the Beamscan measurements are more accurate with more power to make the offsets less significant. Additionally, if this is the case then we can do a fit to the previous data (which was all taken with the same power setting) and simply allow the offset to be a free parameter without affecting the accuracy of the waist calculation. This fit and data coming to an e-log near you soon.

 

Of course, it looks from the plots above (well... the code that produces the plots above) that the waist is actually a little bit small (around 46um) so some adjustment of the last lens back along the beam by about half a cm or so might be required.

 
  4469   Wed Mar 30 20:50:43 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Step 3b: Non-circular? We can fix that...

A quick Beamscan sweep of the beam after the Faraday:

Position A1_13.5%_width A2_13.5%_width

(bench) (um mean) (um mean)

25.8 503.9 478.8

25 477.5 489.0

24 447.1 512.4

21 441.6 604.5

20 476.3 645.4

19 545.4 704.1

18 620.3 762.8

 

After_Faraday.png

 

OK. It looks not too bad - doesn't look too different from what we had. Note that the x axis is in local table units - I found this useful for working out where things were relative to other things (like lenses and the FI) - but it means the beam propagates from right to left in the plot. in other words, the horizontal waist occurs first and is larger than the vertical waist. Also - they're not fitted curves - they're by-eye, best guesses and there's no solution for the vertical that doesn't involve offsets... discussion in a later part of the thread.

 

Anyway! The wonderful thing about this plot is that the horizontal and vertical widths cross and the horizontal focussing at this crossing point is shallower than the vertical. This means that we can put a lens in at the crossing point and rotate it such that the lens is stronger in the horizontal plane. The lens can be rotated until the effective horizontal focal length is right to fix the astigmatism.

 

 

I used a 200mm lens I had handy - a rough check sweeping the Beamscan quickly indicated should be about right though. Adjusting the angle until the beam size at a distant point is approx circular - I then move the profiler and adjust again. Repeat as required. Now… taking some data. with just that lens in:

 

Position A1_13.5%_width A2_13.5%_width

(bench) (um mean) (um mean)

24 371.7 366.1

21 360.3 342.7

20 447.8 427.8

19 552.4 519.0

18 656.4 599.2

17 780.1 709.9

16 885.9 831.1

 

After_Faraday_and_Rotated_Lens.png

 

Well now. That looks quite OK. Fit's a bit rubbish on vertical but looks like a slight offset on the measurement again.

The angle of the lens looks awful, but if it's stupid and it works then it isn't stupid. If necessary, the lens can be tweaked a bit more, but there's always more tweaking possible further down the line and most of the astigmatic behaviour has been removed. It's now just a case of finding a lens that works to give us a 50 um beam at the oven position...

 

 

  4468   Wed Mar 30 20:31:30 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Step 3: Inserting FI and un-eliptical-ification of the beam

The FI set up on it's mount and the beam passes through it - centrally through the apertures on each side. Need to make sure it doesn't clip and also make sure we get 93% through (datasheet specs say this is what we should achieve). We will not achieve this, but anything close should be acceptable.

Setting up for minimum power through the FI is HWP @125deg.

Max is therefore @ 80deg

 

Power before FI = 544 mW

Power after FI =     496 mW (after optimising input polarisation)

Power dumped at input crystal = 8.6mW

Power dumped at input crystal from internal reflections etc = 3.5mW

Power dumped at output crystal on 1st pass = approx 8mW

 

OK. that gives us a 90.625% transmission and a 20.1mW absorption/unexplained loss.

 

Well - OK. The important part about isolators isn't their transmission, it's about how well they isolate. Let's see how much power gets ejected on returning through the isolator…

 

Using a beam splitter to pick off light going into and returning from the FI. A 50/50 BS1-1064-50-1025-45P. And using a mirror near the waist after the FI to send the beam back through. There are better ways to test the isolation performance of FI's but this will suffice for now - really only want to know if there's any reasonable isolation at all or if all of the beam is passing backwards through the device.

 

Power before BS = 536 mW (hmmn - it's gone down a bit)

Power through BS = (can't access ejected on first pass)

Power through FI = 164 mW (BS at odd angle to minimise refractive effect so less power gets through)

Power lost through mirror = 8.3mW (mirror is at normal incidence so a bit transmissive)

 

Using earlier 90.6% measurement as reference, power into FI = 170.83 mW

So BS transmission = 170.83/536 = 0.3187

BS reflectivity therefore = 1 - 0.3187 = 0.6813

 

Power back into FI = Thru FI - Thru mirror = 155.7 mW

 

Power reflected at BS after returning through FI = 2.2mW

Baseline power at BS reflection from assorted internal reflections in FI (blocked return beam) = 1.9mW

Note - these reflections don't appear to be back along the input beam, but they *are* detectable on the power meter.

 

Actual power returning into FI that gets reflected by BS = 0.3 mW

(note that this is in the fluctuating noise level of measurement so treat as an upper limit)

 

Accounting for BS reflectivity at this angle, this gives a return power = 0.3/0.6813 = 0.4403 mW

 

Reduction ratio (extinction ratio) of FI =  0.4403/155.7 = 0.00282

 

Again - note that this upper limit measurement is as rough and ready as it gets. It's easy to optimise this sort of thing later, preferable on a nice open bench with plenty of space and a well-calibrated photodiode. It's just to give an idea that the isolator is actually isolating at all and not spewing light back into the NPRO.

 

Next up… checking the mode-matching again now that the FI is in place. The beam profile was scanned after the FI and the vertical and horizontal waists are different...

  4467   Wed Mar 30 20:14:17 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Additional:

I fired up some old waistplotter routines, and set the input conditions as the measured waist after the lens and used that to work out what the input waist is at the laser. It may not be entirely accurate, but it /will/ be self consistent later on.

 

Vertical waist      = 105.00 um at 6.282 cm after laser output (approx)

Horizontal waist = 144.63 um at 5.842 cm after laser output (approx)

 

Definitely astigmatic.

 

  4466   Wed Mar 30 20:08:34 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Step 2: Getting the beam through the Faraday isolator (FI).

Started out with an f=100mm lens at position 32,T on the bench which gave a decent looking waist of order 100 um in the right sort of position for the FI, but after checking the FI specs, it's limited to 500W/cm^2. In other words, if we have full power from the laser passing into it we'd need a beam width of more than 211 um. Solution? Use an f=150mm lens instead and don't put the FI at the waist. I normally don't put a FI at a waist anyway, for assorted reasons - scattering, thermal lensing, non-linear magnetic fields, the sharp changing of the field components in an area where you want as constant a beam as possible.  Checked with others to make sure they don't do things differently around these parts… Koji says it doesn't matter as long as it passes cleanly through the aperture. So… next step is inserting the Faraday.

The beam profiles in vertical and horizontal around the FI position with the f=150mm lens in place are attached. Note that the FI will be going in at around 0.56m.

Beam_Matching_02c_Vertical.pngBeam_Matching_02c_Horizontal.png

 

 

 

  4465   Wed Mar 30 19:54:19 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

RIght! Overview out of the way - now comes the trivial first bit

 

Step 1: Beam out of the laser - this will be tricky, but we'll see what we can actually measure in this set-up. Can't get the Beamscan head any closer to the laser and using a lambda/2 plate + polariser to control power until the Faraday isolator is in place. Using 1 inch separation holes as reference points for now - need better resolution later, but this is fine for now and gives an idea of where things need to go on the bench. The beam is aligned to the 3rd row up (T) for all measurements, the Beamscan spits out diameters (measuring only the 13.5% values) so convert as required to beam radius and the beam is checked to ensure a reasonable Gaussian profile throughout.

 

Position A1_13.5%_width A2_13.5%_width

(bench) (um mean) (um mean)

32 2166.1 1612.5

31 2283.4 1708.3

30 2416.1 1803.2

29 2547.5 1891.4

27 2860.1 2070.3

26 2930.2 2154.4

25 3074.4 2254.0

24 3207.0 2339.4

 

OK. As expected, this measurement is in the linear region of the beampath - i.e. not close to the  waist position and beyond the Rayleigh length) so it pretty much looks like two straight lines. There's no easy way to get into the path closer to the laser, so reckon we'll just need to infer back from the waist after we get a lens in there. Attached the plot, but about all you really need to get from this is that the beam out of the laser is very astigmatic and that the vertical axis expands faster than the horizontal.

Not terribly exciting, but have to start somewhere.

 

laser_output_non_circular.png

 

 

 

  4464   Wed Mar 30 19:43:33 2011 BryanConfigurationGreen LockingThe wonderful world of mode-matching

Right. I've got a whole load of info and data and assorted musings I've been saving up and cogitating upon before dumping it into these hallowed e-pages. there's so much I'll probably turn it into a threaded entry rather than put everything in one massive page.

An overview of what's coming:

I started out using http://lhocds.ligo-wa.caltech.edu:8000/40m/Advanced_Techniques/Green_Locking?action=AttachFile&do=get&target=modematch_END.png as a reference for roughly what we want to achieve... and from http://nodus.ligo.caltech.edu:8080/40m/100730_093643/efficiency_waist_edit.png we need a waist of about 50um at the green oven. Everything else up to this point is pretty much negotiable and the only defining things that matter are getting the right waist at the doubling oven with enough available power and (after that point) having enough space on the bench to separate off the green beam and match it into the Y arm.

 

So…

Step 1: Measure the properties of the beam out of the laser. Really just need this for reference later because we'll be using more easily measurable points on the bench.


Step 2:
Insert a lens a few cm from the laser to produce a waist of about of a few 100um around the Faraday. Note that there's really quite a lot of freedom here as to where the FI has to be - on the X arm it's around columns 29/30 on the bench, but as long as we get something that works we can get it closer to the laser if we need to.


Step 3:
After inserting the FI need to measure the beam after it (there *will* be some distortion and the beam is non-circular to begin with)


Step 3b:
If beam is non-circular, make it circular.


Step 4:
Insert a lens to produce a 50um waist at the doubling oven position. This is around holes 7/8 on the X arm but again, we're free to change the position of the oven if we find a better solution. The optical set-up is a little bit tight near that side of the bench on the X end so we might want to try aiming for something a bit closer to the middle of the bench? Depends how the lenses work out, but if it fits on the X end it will fit on the Y end.

 
Oh... almost forgot. While I've been doing most of the grunt-work and heavy lifting - thanks go out to Suresh, Kiwamu, Koji, Steve and everyone else who's helped out with discussion of results and assorted assists to numerous to mention.

 

  4463   Wed Mar 30 18:50:57 2011 KojiConfigurationComputer Scripts / ProgramsAdded a sitemap alias

I thought that "m40m" was the traditional alias for the sitemap...

rossa:~>alias
...

m40m ${medm_base} ${medm_newtail} &
...
sitemap medm -x /cvs/cds/rtcds/caltech/c1/medm/sitemap.adl

rossa:~>set|grep medm
medm_base       medm
medm_newtail    -x /opt/rtcds/caltech/c1/medm/sitemap.adl

medm_tail       -x /cvs/cds/caltech/medm/sitemap.adl

Quote:

I added an alias to the sitemap MEDM screen in /cvs/cds/caltech/target/cshrc.40m

Now you can enjoy launching sitemap from a terminal.

alias sitemap 'medm -x /cvs/cds/rtcds/caltech/c1/medm/sitemap.adl'

 

 

  4462   Wed Mar 30 17:01:08 2011 Larisa ThorneUpdateVIDEOCable laying...continued

[Steve, Suresh, Kiwamu, Larisa]

 

Only the PRM/BS cable was laid today.

In one of the previous updates on cable laying, it was noted that the MC2 cable needed an additional 10' and the MC2T needed an additional 15' to reach their destinations.  We cut and put BNC ends on 10' and 15' cables and connected them to the original cables in order to make them long enough.

 

This concludes the laying of new cables. Suresh is currently working on the QUADs...

  4461   Wed Mar 30 16:57:13 2011 kiwamuUpdateGeneralturned off c1aux

[Steve / Kiwamu]

 As a part of the video cable session, we reconnected some power cords on 1Y1 rack.

During the work we momentarily turned off c1aux, which handles DMF, Illumintators, mechanical shutters and the old video epics.

I think it automatically reverted the things, but we may need to check them.

  4460   Wed Mar 30 16:32:29 2011 AidanConfigurationComputer Scripts / ProgramsAdded a sitemap alias

I added an alias to the sitemap MEDM screen in /cvs/cds/caltech/target/cshrc.40m

Now you can enjoy launching sitemap from a terminal.

alias sitemap 'medm -x /cvs/cds/rtcds/caltech/c1/medm/sitemap.adl'

  4459   Wed Mar 30 02:55:02 2011 SureshConfigurationElectronicsRF System : Status and Plans

I have prepared several diagrams outlining the current state of the RF System.

These are uploaded into the svn40m here  and will be kept uptodate as we complete various parts of the task.  These plans have taken into account

the new priorities of the LSC (set out by Koji here )

We (Koji, Kiwamu and I) took stock of the RF cables which we have inherited from the earlier RF system and have made new plans for them.

I took stock of the filters purchased for the modifying the demod boards.  We have pretty much everything we need so I will start modifying the boards right away.   The following table summarises the modifications

 

PD freq # of PDs

LP Filter (U5)

  Demod board

Qty available Inline HP filter Qty available
11 MHz 5 SCLF-10.75 7 - -
22 MHz 1 SCLF-21.4 3 - -
33 MHz 2 SCLF-36 3 SHP-25 1
55 MHz 3 SCLF-65 4 SHP-50 2
110 MHz 1 SCLF-135 3 SHP-100 1
165MHz 3 SCLF-190 1 SHP-150 1

We seem to have a spare SHP-175.  I was wondering where that is supposed to go. 

This is the status and tentative schedule for completing the various tasks.  I have put the dates based on priority and state of the hardware.

 

The RF Cable layout plans are drawn on top of a Lab Layout.  The various subsystems are drawn (not to scale) on separate layers.  The graffle files are located here  .  I thought they might come in handy for others as well.

 

 

  4458   Tue Mar 29 22:29:16 2011 kiwamuUpdateGeneralsome tasks tomorrow

 *  Temporary strain relief for the heliax cables on 1X2 (Steve)

 *  RF diagrams and check lists (Suresh)

      => In the lunch meeting we will discuss the details about what we will do for the RF installation.

 *  Electronics design and plan for Green locking (Aidan / Kiwamu)

      => In the lunch meeting we will discuss the details.

 *  LSC model (Koji)

 *  Video cable session (team)

 * LPF for the laser temperature control (Larisa)

  4457   Tue Mar 29 15:50:21 2011 KojiUpdateElectronicsLow pass filter for X arm laser temperature control

For bode plot:

USE LOG-LOG plot for the amplitude

USE LOG-LINEAR plot for the phase

 

Search "Bode Plot" on web

  4456   Tue Mar 29 15:01:58 2011 Larisa ThorneUpdateElectronicsLow pass filter for X arm laser temperature control

 This is the continuation of http://nodus.ligo.caltech.edu:8080/40m/4402

 

The first picture is of the actual component, where the resistor is 1M and capacitor is 10uF. 

But before the component can be put into place, its transfer function had to be checked to make sure it was doing what we calculated it would do. The results of these are in the graphs generated: frequency vs. gain, and frequency vs. phase.

 

 

 

According to these graphs, we are not achieving the targeted cutoff frequency: need to recalculate and compensate for the extra 100k resistance being encountered.

Attachment 1: DSC_2889.JPG
DSC_2889.JPG
Attachment 2: LPFgraph.pdf
LPFgraph.pdf LPFgraph.pdf LPFgraph.pdf LPFgraph.pdf
  4455   Tue Mar 29 00:00:55 2011 KojiUpdateIOOFixing MC/Freq Divider Box

This is the log of the work on Wednesday 23rd.

1. Power Supply of the freq divider box

Kiwamu claimed that the comparator output of the freq div box only had small output like ~100mV.
The box worked on the electronics bench, we track down the power supply and found the fuse of the +15V line
brew out. It took sometime to notice this fact as the brown-out-LED of the fuse was not on and the power
supply terminal had +15V without the load. But this was because of the facts 1) the fuse is for 24V, and 2)
the large resistor is on the fuse for lighting the LED when the fuse is brown out.

I found another 24V fuse and put it there. Kiwamu is working on getting the correct fuses.

2. MC locking problem

After the hustle of the freq divider, the MC didn't lock. I tracked down the problem on the rack and found
there was no LO for the MC. This was fixed by pushing the power line cable of the AM Stabilizer for the MC LO, which was a bit loose.

  4454   Mon Mar 28 23:51:54 2011 JenneUpdatePSLNew PMC Base Riser Design

Quote:

Its going to need some kind of way to locate the PMC on the top. In the previous design, we had the 3 balls to decouple the body from the base. That design was flawed due to the roughness of the holes in the PMC body.

 Hmmm, so, this was just meant to be a riser that goes underneath the old PMC mount, to raise it from 3" beam height to 4" beam height.  I will make another one that is a complete mount, designed for 4" beam height.  Please hold........... .......... ....... ..... ... .

  4453   Mon Mar 28 22:56:14 2011 ranaUpdatePSLNew PMC Base Riser Design

Its going to need some kind of way to locate the PMC on the top. In the previous design, we had the 3 balls to decouple the body from the base. That design was flawed due to the roughness of the holes in the PMC body.

Also probably need some kind of relief on the bottom. Its possible that it would be OK like this, but I am unsure if the shop can maintain the flatness we want over the whole length and/or the flatness of any given (OLD) optical table over ~8". Its probably not a good idea to have to torque this (aluminum?) to make it conform to the optical table's shape.

  4452   Mon Mar 28 21:12:14 2011 JenneUpdatePSLNew PMC Base Riser Design

I (think) I have finished the new PMC base riser.  The eDrawing of it (so you can view it on any computer) has been uploaded to the PMC wiki page.

I also attach it here, for comments.

Attachment 1: PMC_riser.eprt
  4451   Mon Mar 28 18:22:43 2011 kiwamuUpdateGreen Lockinga mixer school

PLEASE DO NOT DISMANTLE THE SETUP !

Actually we tried looking for a level-3 or a smaller mixer, but we didn't find them at that moment. That's why we kept the level-7 mixer for the coarse path.

As you pointed out we can try an RF amplifier for it.

Quote:

Using 2 dBm for a Level 7 mixer is so bogus, that I will dismantle this as soon as I come over.

http://www.minicircuits.com/products/amplifiers_main.html

 

  4450   Mon Mar 28 18:13:32 2011 ranaUpdateGreen Lockinga mixer school

Using 2 dBm for a Level 7 mixer is so bogus, that I will dismantle this as soon as I come over.

http://www.minicircuits.com/products/amplifiers_main.html

  4449   Mon Mar 28 17:06:15 2011 kiwamuUpdateGreen Lockinga mixer school

In the last week Matt and I modified the MFD configuration because the mixer had been illegally used.

 

mixers.png

Since the output from the comparator is normally about 10 dBm, a 4-way power splitter reduced the power down to 4 dBm in each output port.

In order to reserve a 7 dBm signal to a level-7 mixer, we decided to use an asymmetric power splitter, which is just a combination of 2-way and 3-way splitter shown in the diagram above.

With this configuration we can reserve a 7 dBm signal for a mixer in the fine path.

However on the other hand we sacrificed the coarse path because the power going to the mixer is now 2.2 dBm in each port.

According to the data sheet for the mixer, 1 dB compression point for the RF input is 1dBm. Therefore we put a 1 dB attenuator for the RF port in the coarse system.

In the delay line of the fine path we found that the delay cable was quite lossy and it reduced the power from 2.2 dBm to about 0 dBm.

 

 

  4448   Mon Mar 28 16:24:35 2011 kiwamuUpdateGreen Lockingpower budget on PSL table

   I measured some laser powers associated with the beat-note detection system on the PSL table.

The diagram below is a summary of the measurement. All the data were taken by the Newport power meter.

 The reflection from the beat-note PD is indeed significant as we have seen.

In addition to it the BS has a funny R/T ratio maybe because we are using an unknown BS from the Drever cabinet. I will replace it by a right BS.

RFPD.png

(background)

 During my work for making a noise budget I noticed that we haven't carefully characterize the beat-note detection system.

The final goal of this work is to draw noise curves for all the possible noise sources in one plot.

To draw the shot noise as well as the PD dark noise in the plot, I started collecting the data associated with the beat-note detection system.

 

(Next actions)

 * Estimation and measurement of the shot noise

 * measurement of the PD electrical noise (dark noise)

 * modeling for the PD electrical noise

 * measurement of the doubling efficiency

 * measurement of an amplitude noise coupling in the frequency discriminators

  4447   Mon Mar 28 16:19:23 2011 steveFrogsPhotosvisithing 5th graders

Suresh is captivating his audience with gravity waves on last Friday, March 25

Attachment 1: P1070475.JPG
P1070475.JPG
  4446   Mon Mar 28 15:49:18 2011 josephbUpdateCDSLessons from LST

[Koji,Joe]

PART 1:

Koji was unable to build his c1lst model first thing this morning.  Turns out there was  a bug with RCG parser that was introduced on Friday when we did the RCG updates.  We talked Alex who did a quick comment fix.  The diff is as follows:

Index: Parser3.pm
==============================

=====================================
--- Parser3.pm  (revision 2328)
+++ Parser3.pm  (working copy)
@@ -1124,8 +1124,8 @@
  print "Flattening the model\n";
  flatten_nested_subsystems($
root);
  print "Finished flattening the model\n";
-  CDS::Tree::do_on_nodes($root, \&remove_tags, 0, $root);
-  print "Removed Tags\n";
+  #CDS::Tree::do_on_nodes($root, \&remove_tags, 0, $root);
+  #print "Removed Tags\n";
  #print "TREE\n";
  #CDS::Tree::print_tree($root);
  CDS::Tree::do_on_nodes($root, \&remove_busses, 0, $root);

This was some code to remove TAGs from the .mdl file for some reason which I do not understand at this time.  I will ask tommorrow in person so I can understand the full story.

PART 2:

Koji then rebuilt and started the c1lst process.  This is his new test version of the LSC code.  We descovered (again) that when you activate too many DAQ channels (simply uncommenting them, not even recording them with activate=1 in the .ini file) that the frame builder crashes.  In addition, the c1lsc machine, which the code was running on, also hard crashed.

When a channel gets added to the .ini file (or uncommented) it is sent to the framebuilder, irregardless of whether its recorded or not by the frame builder.  There is only about 2 megabytes per second bandwidth per computer.  In this case we were trying to do something like 200 channels * 16384 Hz * 4 bytes = 13 megabytes per second.

The maximium number of 16384 channels is roughly 30, with little to no room for anything else.  In addition, test points use the same allocated memory structure, so that if you use up all the capacity with channels, you won't be able to use testpoints to that computer (or thats what Alex has led me to believe).

The daqd process then core dumped and was causing all sorts of martian network slowdowns.  At the same time, the c1lsc computer crashed hard, and all of the front end processes except for the IOP on c1sus crashed.

We rebooted c1lsc, and restarted the c1sus processes using the startc1SYS scripts.  However, the c1susfe.ko apparently got stuck in a wierd state.  We were completely unable to damp the optics and were in general ringing them up severely.  We tried debugging, including several burt restores and single path checks.

Eventually we decided to reboot the c1sus machine after a bit of debugging.  After doing a burt restore after the reboot, everything started to damp and work happily.  My best guess is the kernel module crashed in a bad way and remained in memory when we simply did the restart scripts.

 

  4445   Mon Mar 28 15:18:04 2011 josephbUpdateCDSCDS updates on Friday

Last Friday, we discovered a bug in the RCG where the delay part was not actually delaying.  We reported this to Alex who promptly put a fix in the same day.  This allowed Matt's newly proposed frequency discriminator to work properly.

It also required a checkout of the latest RCG code (revision 2328), and rebuild of the various codes.  We backed up all the kernel and executables first such as mbuf.ko and awgtpman.

We did the following:

1) Log into the fb machine.

2) Go to /opt/rtcds/caltech/c1/core/advLigoRTS/src/drv/mbuf and run make.  Copy the newly built mbuf.ko file to /diskless/root/modules/2.6.34.1/kernel/drivers/mbuf/mbuf.ko on the fb machine.

3) Use "sudo cp" to copy the newly built mbuf.ko file to /diskless/root/modules/2.6.34.1/kernel/drivers/mbuf/

4) Go to /cvs/cds/rtcds/caltech/c1/core/advLigoRTS/src/gds and run make.

5) Copy the newly built awgtpman executable to /opt/rtcds/caltech/c1/target/gds/bin/

6) Go to /opt/rtcds/caltech/c1/core/advLigoRTS/src/mx_stream/ and run make.

7) Copy the newly built mx_stream executable to /opt/rtcds/caltech/c1/target/fb/

  4444   Fri Mar 25 11:16:19 2011 josephbFrogsGreen Lockingdigital frequency counting

I modified the c1gfd.mdl simulink model.  I made a backup as c1gfd_20110325.mdl. 

The first change was to use a top_names block to put everything in.  The block is labeled ALS.  So all the channels will now be C1:ALS-GFD_SOMETHING.  This means medm channel names will need to be updated.  Also, the filter modules need to be updated in foton because of this.

I then proceeded to add the suggested changes made by Matt.  To avoid a divide by zero case, I added a saturation part which saturates at 1e-9 (note this is positive) and 1e9.

Quote:

Today we tried the Schmitt trigger DFD, and while it works it does not improve the noise performance.  At least part of our problem is coming from the discrete nature of our DFD algorithm, so I would propose that an industrious day job person codes up a new DFD which avoids switching.  We can probably do this by mixing the input signal (after high-passing) with a time-delayed copy of itself... as we do now, but without the comparator.  This has the disadvantage of giving an amplitude dependent output, but since we are working in the digital land we can DIVIDE.  If we mix the signal with itself (without delay) to get a rectified version, and low-pass it a little, we can use this for normalization.  The net result should be something like:

output = LP2[ s(t) * s(t - dt) / LP1[ s(t) * s(t) ]],

where s(t) is the high-passed input and LP is a low-pass filter.  Remember not to divide by zero.

 

 

Attachment 1: C1GFD.png
C1GFD.png
  4443   Fri Mar 25 08:58:38 2011 AidanUpdateTreasureCleared stuff off SP table

I tidied up some of the stuff that was on the SP table. The ISS box that has been sitting on there for months is now underneath the X-arm on top of the spare Marconi which is stored there.

 

 

Attachment 1: ISS_box.JPG
ISS_box.JPG
  4442   Fri Mar 25 01:27:29 2011 mevansFrogsGreen Lockingdigital frequency counting

Today we tried the Schmitt trigger DFD, and while it works it does not improve the noise performance.  At least part of our problem is coming from the discrete nature of our DFD algorithm, so I would propose that an industrious day job person codes up a new DFD which avoids switching.  We can probably do this by mixing the input signal (after high-passing) with a time-delayed copy of itself... as we do now, but without the comparator.  This has the disadvantage of giving an amplitude dependent output, but since we are working in the digital land we can DIVIDE.  If we mix the signal with itself (without delay) to get a rectified version, and low-pass it a little, we can use this for normalization.  The net result should be something like:

output = LP2[ s(t) * s(t - dt) / LP1[ s(t) * s(t) ]],

where s(t) is the high-passed input and LP is a low-pass filter.  Remember not to divide by zero.

 

  4441   Thu Mar 24 19:48:13 2011 Aidan, KiwamuUpdateGreen LockingDesigns for permanent electronics for ALS

Kiwamu and I looked at all the electronics that are currently in place for the green locking on the X-arm and have made a set of block diagrams of the rack mounted units that we should build to replace the existing ... "works of art" that sprawl around out there at the moment.

Main items

1. "ETM Green Oscillator/PDH support box". Not a great name but this would provide the local oscillator signal for the end PDH (with a controllable phase rotator) as well as the drive oscillator for the end laser PZT. Since we need to hit a frequency of 216.075kHz with a precision that Kiwamu needs to determine, we'd need to be able to tune the oscillator ... it needs to be a VCO. It'd be nice to be able to measure the output frequency so I've suggested dividing it down by N times to put it into the DAQ - maybe N = 2^7 = 128x to give a measured frequency of around 1.7kHz. Additionally this unit will sum the PDH control signal into the oscillation. This box would support the Universal PDH box that is currently at the X-end.

2. "Vertex X-arm beatnote box" - this basically takes the RF and DC signals from the beatnote PD and amplifies them. It provides a monitor for the RF signal and then converts the RF signal into a square wave in the comparator.

3. "Mixer Frequency Discriminator" - just the standard MFD setup stored in a box. For temperature stability reasons, we want to be careful about where we store this box and what it is made of. That's also the reason that this stage is separated from the X-arm beatnote box with it's high-power amps.

Other things

4. RS232 and EPICS control of the doubling ovens

5. Intensity stabilization of the End Laser

P.S. I used Google Diagrams for the pictures.

Attachment 1: GreenLockingElectronics.pdf
GreenLockingElectronics.pdf
Attachment 2: GreenEndPDHsupportboxandLO.pdf
GreenEndPDHsupportboxandLO.pdf
Attachment 3: VertexBeatnoteAmplifierandComparator.pdf
VertexBeatnoteAmplifierandComparator.pdf
Attachment 4: MixerFrequencyDiscriminator.pdf
MixerFrequencyDiscriminator.pdf
  4440   Thu Mar 24 16:33:32 2011 BryanConfigurationGreen LockingPSL vs Y arm laser temperature pairing

X_arm and Y_arm vs PSL comparison.

 

 

Just a quick check of the performance of the X arm and Y arm lasers in comparison to the PSL. Plotting the data from the X arm vs PSL and Y arm vs PSL on the same plot shows that the X arm vs the PSL has no observable trending of mode-hopping in the laser, while the Y arm vs the PSL does. Suspect this is due to the fact that the X arm and PSL are both Innolight lasers with essentially identical geometry and crystals and they'll tend to mode-hop at roughly the same temperatures - note that the Xarm data is rough grained resolution so it's likely that any mode-hop transitions have been skipped over. The Lightwave on the other hand is a very different beast and has a different response, so won't hop modes at the same temperatures.

Given how close the PSL is to one of the mode-transition regions where it's currently operating (32.26 degC) it might be worth considering shifting the operating temperature down one degree or so to around 31 degC? Just to give a bit more headroom. Certainly worth bearing in mind if problems are noticed in the future.

PSL_vs_X-end_T_data_fine.jpg

  4439   Thu Mar 24 15:30:59 2011 BryanConfigurationGreen LockingPSL vs Y arm laser temperature pairing
Fine-grained temperature vs temperature data around the current operating point of the PSL laser.
 
The last set of data was taken in 1 degreeC steps, but we want a bit more detail to find out what happens around the current PSL operating point. So we took some data with a 0.1 degC resolution.

The good news is that we seem to be running in a linear region of the PSL laser with a degree or so of range before the PSL Innolight laser starts to run multi-mode. On the attached graph we are currently running the PSL at 32.26degrees (measured) which puts us in the lower left corner of the plot. The blue data is the Lightwave set temperature (taken from the display on the laser controller) and the red data is the Lightwave laser crystal measured temperature (taken from the 10V/degC calibrated diagnostic output on the back of the laser controller - between pins 2 and 4).

The other good news is that we can see the transition between the PSL laser running in one mode and running in the next mode along. The transition region has no data points because the PMC has trouble locking on the multi-mode laser output - you can tell when this is happening because, as we approach the transition the PMC transmitted power starts to drop off and comes back up again once we're into the next mode region (top left portion of the plot).

 

The fitted lines for the region we're operating in are:

Y_arm_Temp_meas = 0.95152*T_PSL + 3.8672

Y_arm_Temp_set = 0.87326*T_PSL + 6.9825

Temp_Beating_Day02_html_m2d719182.jpg

 

  4438   Thu Mar 24 13:56:05 2011 josephbUpdateelogelog restarted at 1:55pm

Restarted elog.

  4437   Thu Mar 24 13:50:30 2011 BryanConfigurationGreen LockingY arm laser

 Just a quick update... the Lightwave laser has now been moved up to the end of the Y arm. It's also been mounted on the new mounting block and heatsinks attached with indium as the heat transfer medium.

A couple of nice piccies...IMG_0188.JPG

Attachment 2: IMG_0190.JPG
IMG_0190.JPG
  4436   Thu Mar 24 01:16:19 2011 SureshSummaryGreen LockingY-END green equipment is all available
There was a 2" mirror mount among the spares on the PSL table.  It has a window LW-3-2050 UV mounted in it.  I
have moved it to the Y-end table.  We seem to have run out of 2" mirror mounts ...
  4435   Wed Mar 23 19:16:17 2011 AidanSummaryGreen LockingY-END green equipment is all available

With the exception of a 2" mirror mount, I've confirmed that we have everything for the Y-end green production and mode-matching.

We need to calculate a mode-matching solution for the Lightwave laser so that it gives the correct beam size in the doubling crystal.

Additionally, Rana has suggested that we change the pedestals from the normal 1" diameter pedestal+fork combo  to the 3/4" diameter posts and wider bases that are used on the PSL table (as shown in the attached image).

Attachment 1: three-quarter_inch_pedestal.jpg
three-quarter_inch_pedestal.jpg
  4434   Wed Mar 23 16:06:20 2011 Larisa ThorneUpdateElectronicsUpdate on cable laying

 [Steve, Suresh, Larisa]

The following cables were laid today: ETMYT, ETMY, IFOPO, MC1, OMCR, AS Spare, and MC2T.

 

Though the paper suggested 135' for the MC2T, we used a 110'. This is too short: need at least another 15' for the MC2T.

The RCR cable wasn't crossed off on the list, but a cable exists at the RCR cable which is black and is labeled (old label, 75 ohms)

There was no indication of which length was needed for MC1, so a 95' cable was used.

ELOG V3.1.3-